OpenStudy (anonymous):

heres another one, write an equation for the line that is perpendicular to the given line and that contains the given point. 3x + 2y = -10; (-9,-2)

OpenStudy (anonymous):

Perpendicular is a different story. If the orginal line have the slope of 3. Then a new perpendicular line must have a slope of the negative reciprocal. This means that the slope must be -(1/3).

OpenStudy (anonymous):

What is the slope of the line 3x + 2y = -10? If you put it into slope-intercept form, you'll see that y = (-3/2) x - 5. So the slope is -3/2. The slope of a perpendicular line to 3x + 2y = -10 must then have slope 2/3 (the negative reciprocal of -3/2). So then we want this perpendicular line to pass through (-9, -2). So we can use “point-slope” form of a line: y – y1 = m*(x – x1). So y – (-2) = (2/3) * (x – (-9)). Simplify to get the equation of the line which is perpendicular to your first line and going through the point (-9, -2).